# Studying propagating turbulent structures in the near wake of a sphere using Hilbert proper orthogonal decomposition

### v1

@@ 4.1 POD and HPOD Modes
The first HPOD mode, the real and imaginary parts of which are shown in figure 4a and 4b, respectively, is qualitatively similar in structure to the first two POD modes, shown in figure 3a and 3b, respectively. The streamwise components of the first HPOD mode, shown in figures 4a(i) and 4b(i) for the real and imaginary parts, respectively, consist of alternating positive and negative regions in the streamwise direction and is antisymmetric about the streamwise axis. The transverse components, shown in figures 4a(ii) and 4b(ii) for the real and imaginary parts, respectively, consist of symmetric regions of alternating sign. This structure is very similar to that of the first two POD modes, with the exception of the ends of the measurement domain. This difference is due to the filtering of the Hilbert transform, which places more emphasis on the components of the structure which propagate in the streamwise direction. The effect of this filtering is seen most clearly comparing figures 4b(i) and 3b(i) around $x/D=5$. The second HPOD mode features the same V-shaped structure in the streamwise components, the real and imaginary parts of which are shown in figures 4c(i) and 4d(i), respectively, present in the fourth and fifth POD modes, shown in figures 3d(i) and 3e(i), respectively. The transverse components of the real and imaginary parts of the second HPOD mode are antisymmetric with regions of alternating sign in the streamwise direction, as shown in figures 4c(ii) and 4d(ii), respectively. While similar to transverse components of the fourth and fifth POD modes, shown in figures 3d(ii) and 3e(ii), respectively, the HPOD modes are significantly more symmetrical/anti-symmetrical than the POD modes, due to the filtering effect of the Hilbert transform.
@@
The third HPOD mode, the real and imaginary components of which are shown in figures 4e and 4f, have a similar structure to the sixth and eight POD modes, shown in figures 3f and 3h, respectively. Namely, antisymmetric streamwise velocity regions and symmetric transverse velocity regions of alternating sign in the streamwise direction, with a shorter wavelength than the first HPOD mode and the first two POD modes. The fourth HPOD mode, the real and imaginary components of which are shown in figures 4g and 4h, have a similar structure to the ninth and tenth POD modes, shown in figures 3i and 3j, respectively. The streamwise components of these modes feature symmetrical ovular regions of alternating sign in the streamwise direction in the near wake becoming V-shaped structures around $x/D=2$, while the transverse components feature antisymmetric regions of alternating sign in the streamwise direction.
@@ 4.2 Pairing POD Modes using HPOD Modes
As $\boldsymbol{f}(\boldsymbol{x},t)$ is given by the real part of $\boldsymbol{f}^{a}(\boldsymbol{x},t)$, it can be expressed in terms of the HPOD modes as
$\boldsymbol{f}(\boldsymbol{x},t)=\Re\big{(}\boldsymbol{f}^{a}(\boldsymbol{x},t)\big{)}=\sum_{k=1}^{K}\Re\big{(}a_{k}^{a}(t)\psi^{a}_{k}(\boldsymbol{x})\big{)},$ (36)
@@
$\boldsymbol{f}(\boldsymbol{x},t)=\Re\big{(}\boldsymbol{f}^{a}(\boldsymbol{x},t)\big{)}=\sum_{k=1}^{K}\Re\big{(}a_{k}^{a}(t)\big{)}\Re\big{(}\psi^{a}_{k}(\boldsymbol{x})\big{)}-\Im\big{(}a_{k}^{a}(t)\big{)}\Im\big{(}\psi_{k}^{a}(\boldsymbol{x})\big{)}.$ (37)
As the real and imaginary parts of $a_{k}^{a}(t)$ can be expressed as $|a^{a}_{k}(t)|\cos(\phi^{a}_{k}(t))$ and $-|a^{a}_{k}(t)|\sin(\phi^{a}_{k}(t))$, respectively, the shape of the $k^{th}$ HPOD mode at the phase angle $\phi_{k}^{a}(t)$ is given by
$\frac{\psi_{k}^{a}(\boldsymbol{x},\phi_{k}^{a}(t))}{|a_{k}^{a}(t)|}=\cos\big{(}\phi_{k}^{a}(t)\big{)}\Re\big{(}\psi_{k}^{a}(\boldsymbol{x})\big{)}+\sin\big{(}\phi_{k}^{a}(t)\big{)}\Im\big{(}\psi_{k}^{a}(\boldsymbol{x})\big{)}.$ (38)
@@
where $\Delta\phi_{k,j}^{a}$ is the phase offset of the $k^{th}$ HPOD mode which results in the best match of the $j^{th}$ POD mode. The negative of the imaginary part is used to account for the negative introduced by the Hilbert transform for positive instantaneous frequencies, given in equation (16), which results in the phase of the HPOD modes being reversed relative to the motion of the structure.
The phase angles at which the $k^{th}$ HPOD mode matches each of the POD modes can be determined simultaneously by
@@
The correlation coefficient of the $k^{th}$ HPOD mode at $\Delta\phi^{a}_{k,j}$ and the $j^{th}$ POD mode is
$\mathcal{R}_{k,j}=\frac{\operatorname{cov}\big{(}\psi_{k}^{a}(\boldsymbol{x},\Delta\phi^{a}_{k,j}),\psi_{j}(\boldsymbol{x})\big{)}}{\sqrt{\operatorname{var}\big{(}\psi_{k}^{a}(\boldsymbol{x},\Delta\phi^{a}_{k,j})\big{)}\operatorname{var}\big{(}\psi_{j}(\boldsymbol{x},)\big{)}}},$ (42)
and the corresponding $R^{2}$ value is
$R^{2}_{k,j}=1-\frac{\sum\big{(}\psi_{k}^{a}(\boldsymbol{x},\Delta\phi^{a}_{k,j})-\psi_{j}(\boldsymbol{x})\big{)}^{2}}{\sum\big{(}\psi_{j}(\boldsymbol{x})\big{)}^{2}}.$ (43)
@@
The HPOD modes which best match each of the first ten POD modes, and the corresponding phase offset, correlation coefficient, and $R^{2}$ value, are shown in table 3.
POD Mode, $j$ HPOD Mode, $k$ $\Delta\phi_{k,j}^{a}$ $\mathcal{R}_{k,j}$ $R^{2}_{k,j}$
1 1 $1.71\pi$ 0.89 0.78
2 1 $1.18\pi$ 0.97 0.93
3 3 $0.18\pi$ 0.62 0.26
4 2 $1.70\pi$ 0.84 0.65
5 2 $1.21\pi$ 0.96 0.91
6 3 $1.52\pi$ 0.82 0.65
7 4 $1.70\pi$ 0.63 0.29
8 3 $0.88\pi$ 0.76 0.50
9 4 $1.10\pi$ 0.85 0.70
10 4 $0.50\pi$ 0.80 0.49
@@
Table 3: HPOD modes and phase shifts for best match of the first ten POD modes, and the corresponding $\mathcal{R}_{k,j}$ values.
The first and second POD modes are the best matches to the first HPOD mode, with correlation coefficients of 0.89 and 0.97, and $R^{2}$ values of 0.78 and 0.93, respectively. The corresponding phase angles are $1.71\pi$ and $1.18\pi$, which are approximately $90^{\circ}$ apart. This is consistent with the separation of the real and imaginary parts of the HPOD modes, suggesting that the relationship between these POD modes is similar to that of the real and imaginary parts of the first HPOD mode. Similarly, the fourth and fifth POD modes are the best matches to the second HPOD mode, with correlation coefficients of 0.84 and 0.96, and $R^{2}$ values of 0.65 and 0.91, respectively. The corresponding phase angles are $1.7\pi$ and $1.21\pi$, which are also separated by approximately $90^{\circ}$.
The sixth and eighth POD modes are the best match to the third HPOD mode, with correlation coefficients of 0.82 and 0.76, and $R^{2}$ values of 0.65 and 0.50, respectively. The separation of the corresponding phase angles, $1.52\pi$ and $0.88\pi$, respectively, is greater than $90^{\circ}$. This is also the case for the ninth and tenth POD modes, which are the best matches to the fourth HPOD mode at phase angles of $1.10\pi$ and $0.50\pi$, respectively. The corresponding correlation coefficients and $R^{2}$ values of these POD modes are 0.85 and 0.80, and 0.70 and 0.49, respectively.
The phase average of the paired POD modes were compared to the corresponding HPOD modes at phase angles from 0 to $2\pi$ in increments of $0.01\pi$, accounting for the phase angle shift in the first POD mode in each pair, $\Delta\phi_{k,i}$. The $R^{2}$ values of each phase average
$R_{k,(i,j)}^{2}=1-\frac{\sum{\Big{(}\psi^{a}_{k}(\boldsymbol{x},\phi+\Delta\phi^{a}_{k,i})-\widetilde{\psi_{i,j}}(\boldsymbol{x},\phi)\Big{)}}^{2}}{\sum{\big{(}\widetilde{\psi_{i,j}}(\boldsymbol{x},\phi)\big{)}}^{2}},$ (44)
was calculated for each case. The joint probability density functions of the phase angle of the HPOD modes and the phase angle of the corresponding pairs of POD modes, adjusted for the phase angle shift between the HPOD mode and the first POD mode of each pair.
HPOD Mode $k$ POD Modes $i,j$ $R_{k,(i,j)}^{2}$
1 1,2 0.81
2 4,5 0.76
3 6,8 0.57
4 9,10 0.61
@@
Table 4: $R_{k,(i,j)}^{2}$ values of the phase-averaged HPOD modes and the corresponding POD modes.
@@
Figure 5: Phase-averaged (a) streamwise and (c) transverse components of the first and second POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the first HPOD mode, adjusted for the phase angle offset $\Delta\phi_{1,1}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
The streamwise component of the phase-averaged first and second POD modes, shown on figure 5a, is good agreement with that of the phase average of the first HPOD mode at the adjusted phase angle, shown in figure 5b, with the main differences in the structures being a result of the filtering from the Hilbert transform. Namely, the intensity of the structures in the POD mode is higher directly behind the sphere than the HPOD modes, while the intensity in the structures further downstream is higher in the HPOD modes. This difference is also seen in the transverse comments, shown in figures 5c and 5d for the POD modes and HPOD modes, respectively.
@@
Additionally, the structures in the near wake of the sphere in the POD modes, shown in figure 5a(iv), are filtered out of the HPOD mode, shown in figure 5b(iv), as they do not propagate downstream. Despite these differences, it is clear that the first and second POD modes and the first HPOD mode represent the same propagating structure consisting of planar antisymmetric streamwise fluctuations and alternating positive and negative transverse fluctuations. This structure is consistent with flapping of the wake, and the similarity of the third POD mode, shown in figure 3c, to the first and second suggests that the flapping also has an out-of-plane component.
@@
Figure 6: Phase-averaged (a) streamwise and (c) transverse components of the fourth and fifth POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the second HPOD mode, adjusted for the phase angle offset $\Delta\phi_{2,4}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
@@
While the streamwise component of the phase-averaged fourth and fifth POD modes, shown in figure 6a, is similar to that of the second HPOD mode, shown in figure 6b, the latter demonstrates significantly stronger symmetry, due to the inherent filtering of propagating features by HPOD. This is most clearly seen when comparing figures 6a(ii) and 6a(vi) to figures 6b(ii) and 6b(vi), respectively, where the V-shaped structures in the HPOD case remain symmetric downstream, while they become asymmetric in the case of the POD modes. The transverse component of the second HPOD mode, shown in figure 6d, shows a stronger antisymmetry than the transverse components of the phase average of the fourth and fifth POD modes, shown in figure 6c.
The variation between the phase average of the fourth and fifth POD modes and the second HPOD mode highlights the filtering effect of HPOD, which prioritises the structures that are periodic in the direction of the Hilbert transform. While the fourth and fifth POD modes are highly asymmetric, particularly after $x/D>3$, the second HPOD mode is symmetric in the streamwise component and antisymmetric in the transverse direction. The asymmetry in the POD modes is likely due to the three-dimensionality in the flow being more significant in the structure represented by the fourth and fifth POD modes than that represented by the first and second POD modes. In this case, the use of HPOD is beneficial for isolating the propagating structures in the flow in the presence of asymmetry in the POD modes. The alternating positive and negative streamwise fluctuations in these modes suggests that they represent a pulsating motion in the wake of the sphere.
@@
Figure 7: Phase-averaged (a) streamwise and (c) transverse components of the sixth and eighth POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the third HPOD mode, adjusted for the phase angle offset $\Delta\phi_{3,6}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
@@
The streamwise component of the phase average of the sixth and eight POD modes, shown in figure 7a, is qualitatively similar to that of the third HPOD mode, shown in figure 7b. However, the alternating positive and negative structures are better defined in the HPOD mode than the phase average of the POD modes. This is most evident when comparing figures 7a(i) and 7a(ii) to figures 7b(i) and 7b(ii), in which the positive region in the HPOD modes around $x/D=3$ and $y/D>0$ is not present in the phase-averaged POD modes. This is most likely due to the structure moving in the out-of-plane direction as it propagates, which is filtered out by the Hilbert transform. Additionally, the intensity of the structures is generally lower in the HPOD modes than those in the POD modes due to the filtering of the non-propagating components.
The transverse component of the phase average of the sixth and eighth POD modes, shown in figure 7c, shows a transverse shifting of the structures of alternating sign, particularly in figures 7c(i)-(iii). This shifting is not present in the transverse component of the third HPOD mode, shown in figure 7d. The shifting of the structures in the transverse velocity of the POD modes is consistent with the the three-dimensional motion of the wake, and coupled with the behaviour of the streamwise components of the POD modes. The structure represented by these modes appears to be a flapping in the wake with a shorter wavelength than that represented by the first and second POD modes and the first HPOD mode.

### v2

@@ 4.1 POD and HPOD Modes
The streamwise component of the second HPOD mode, the real and imaginary parts of which are shown in figures 4c(i) and 4d(i), respectively, features the same V-shaped structure present in the fourth and fifth POD modes, shown in figures 3d(i) and 3e(i), respectively. The transverse components of the real and imaginary parts of the second HPOD mode are antisymmetric with regions of alternating sign along the streamwise direction, as shown in figures 4c(ii) and 4d(ii), respectively. While similar to the transverse components of the fourth and fifth POD modes, shown in figures 3d(ii) and 3e(ii), respectively, the HPOD modes are significantly more symmetric and antisymmetric across the centreline than the POD modes, due to the accentuation of propagating structures in the HPOD modes.
@@
The third HPOD mode, the real and imaginary components of which are shown in figures 4e and 4f, respectively, has a similar structure to the sixth and eighth POD modes, shown in figures 3f and 3h, respectively. Specifically, it exhibits antisymmetric streamwise velocity regions and symmetric transverse velocity regions of alternating sign along the streamwise direction, with a shorter wavelength than the first HPOD mode and the first two POD modes. The fourth HPOD mode, the real and imaginary components of which are shown in figures 4g and 4h, has a similar structure to the ninth and tenth POD modes, shown in figures 3i and 3j, respectively. The streamwise component of this mode features symmetric ovular regions of alternating sign in the streamwise direction in the near wake, becoming V-shaped structures around $x/D=2$, while the transverse component features antisymmetric regions of alternating sign in the streamwise direction.
@@ 4.2 Pairing POD Modes using HPOD Modes
As $\boldsymbol{f}(\boldsymbol{x},t)$ is given by the real part of $\boldsymbol{f}^{a}(\boldsymbol{x},t)$, it can be expressed using the HPOD modes as
$\boldsymbol{f}(\boldsymbol{x},t)=\Re\big(\boldsymbol{f}^{a}(\boldsymbol{x},t)\big)=\sum_{k=1}^{K}\Re\big(a^{a}_{k}(t)\psi^{a}_{k}(\boldsymbol{x})\big),$ (37)
@@
$\boldsymbol{f}(\boldsymbol{x},t)=\Re\big(\boldsymbol{f}^{a}(\boldsymbol{x},t)\big)=\sum_{k=1}^{K}\Re\big(a^{a}_{k}(t)\big)\Re\big(\psi^{a}_{k}(\boldsymbol{x})\big)-\Im\big(a^{a}_{k}(t)\big)\Im\big(\psi^{a}_{k}(\boldsymbol{x})\big).$ (38)
As the real and imaginary parts of $a^{a}_{k}(t)$ can be expressed as $|a^{a}_{k}(t)|\cos(\phi^{a}_{k}(t))$ and $|a^{a}_{k}(t)|\sin(\phi^{a}_{k}(t))$, respectively, the shape of the $k^{th}$ HPOD mode for the phase angle $\phi^{a}_{k}(t)$ is given by
$\frac{\psi^{a}_{k}(\boldsymbol{x},\phi^{a}_{k}(t))}{|a^{a}_{k}(t)|}=\cos\big(\phi^{a}_{k}(t)\big)\Re\big(\psi^{a}_{k}(\boldsymbol{x})\big)+\sin\big(\phi^{a}_{k}(t)\big)\Im\big(\psi^{a}_{k}(\boldsymbol{x})\big).$ (39)
@@
where $\Delta\phi_{k,j}^{a}$ is the phase offset of the $k^{th}$ HPOD mode which provides the best match to the $j^{th}$ POD mode. The negative sign on the imaginary part accounts for the negative introduced by the Hilbert transform associated with positive instantaneous frequencies, given in equation (16), which accounts for the phase of the HPOD modes being reversed relative to the motion of the structure.
The phase angles at which the $k^{th}$ HPOD mode matches each of the POD modes can be determined by solving
@@
The correlation coefficient between the $k^{th}$ HPOD mode at $\Delta\phi^{a}_{k,j}$ and the $j^{th}$ POD mode is given by
$\mathcal{R}_{k,j}(\Delta\phi^{a}_{k,j})=\frac{\operatorname{cov}\big(\psi^{a}_{k}(\boldsymbol{x},\Delta\phi^{a}_{k,j}),\psi_{j}(\boldsymbol{x})\big)}{\sqrt{\operatorname{var}\big(\psi^{a}_{k}(\boldsymbol{x},\Delta\phi^{a}_{k,j})\big)\operatorname{var}\big(\psi_{j}(\boldsymbol{x})\big)}},$ (43)
and the corresponding $R^{2}_{k,j}$ value is
$R^{2}_{k,j}(\Delta\phi^{a}_{k,j})=1-\frac{\sum_{\boldsymbol{x}}{\Big(\psi^{a}_{k}(\boldsymbol{x},\Delta\phi_{k,j}^{a})-\psi_{j}(\boldsymbol{x})\Big)}^{2}}{\sum_{\boldsymbol{x}}{\Big(\psi_{j}(\boldsymbol{x})\Big)}^{2}},$ (44)
@@
where $\boldsymbol{x}$ represents the spatial coordinates of the modes. The $\mathcal{R}_{k,j}$ and $R_{k,j}^{2}$ values were calculated for the combined streamwise and transverse components of the modes, as well as individually for each component, denoted by superscripts u and v for the streamwise and transverse components, respectively. The HPOD modes which best match each of the first ten POD modes, and the corresponding phase offset $\Delta\phi^{a}_{k,j}$, correlation coefficient $\mathcal{R}_{k,j}$, and $R_{k,j}^{2}$ value, are shown in table 3.
POD Mode, $j$ HPOD Mode, $k$ $\Delta\phi_{k,j}^{a}$ $\mathcal{R}_{k,j}$ $\mathcal{R}_{k,j}^{u}$ $\mathcal{R}_{k,j}^{v}$ $R^{2}_{k,j}$ $(R^{2}_{k,j})^{u}$ $(R^{2}_{k,j})^{v}$
1 1 $1.64\pi$ 0.88 0.86 0.90 0.74 0.71 0.75
2 1 $1.15\pi$ 0.98 0.96 0.99 0.95 0.92 0.96
3 3 $0.65\pi$ 0.57 0.46 0.66 0.28 0.10 0.38
4 2 $1.71\pi$ 0.86 0.85 0.88 0.68 0.70 0.74
5 2 $1.21\pi$ 0.96 0.95 0.97 0.84 0.87 0.83
6 3 $1.49\pi$ 0.87 0.83 0.92 0.75 0.68 0.80
7 4 $0.21\pi$ 0.56 0.59 0.56 0.15 0.20 0.27
8 3 $0.92\pi$ 0.80 0.65 0.90 0.60 0.35 0.78
9 4 $1.67\pi$ 0.86 0.84 0.91 0.73 0.63 0.83
10 4 $1.16\pi$ 0.86 0.86 0.92 0.68 0.57 0.85
@@
Table 3: HPOD modes and phase shifts for the best match of the first ten POD modes, and the corresponding $\mathcal{R}_{k,j}$ values. Superscripts u and v denote the streamwise and transverse components, respectively.
The correlation coefficient between the $j^{th}$ POD mode and the $k^{th}$ HPOD mode as a function of streamwise position is given by
$\mathcal{R}_{k,j}(x,\Delta\phi^{a}_{k,j})=\frac{\operatorname{cov}\big(\psi^{a}_{k}(x,\Delta\phi^{a}_{k,j}),\psi_{j}(x)\big)}{\sqrt{\operatorname{var}\big(\psi^{a}_{k}(x,\Delta\phi^{a}_{k,j})\big)\operatorname{var}\big(\psi_{j}(x)\big)}},$ (45)
and are shown for the matched HPOD modes and POD modes in figure 5.
Figure 5: Correlation coefficient between each paired POD mode and the corresponding HPOD mode at the best match phase angle as a function of streamwise position. Superscripts u and v denote the streamwise and transverse components, respectively.
The first and second POD modes exhibit the strongest correspondence to the first HPOD mode, with correlation coefficients of 0.88 and 0.98, and $R_{k,j}^{2}$ values of 0.74 and 0.95, respectively. The corresponding phase angles are $1.64\pi$ and $1.15\pi$, which are approximately $\pi/2$ apart and consistent with the separation of the real and imaginary parts of the HPOD modes. This suggests that the relationship between these POD modes mirrors that of the real and imaginary components of the first HPOD mode. The correlation coefficient between these POD and HPOD modes is consistently close to unity, with troughs in the correlation occurring due to the values of the modes approaching zero at $x/D=2.6$ and $x/D=4.4$, as shown in figure 5a. Additional reductions occur near the domain boundaries, with the trough at the upstream end being significantly deeper. While the correlation of the combined and streamwise components decreases to approximately 0.5, the correlation of the transverse component approaches -1. This indicates that the transverse components of the first POD mode and the first HPOD mode have opposite signs close to the sphere. This discrepancy likely results from a combination of spectral leakage at the domain edge and the difference in length of the structures in the HPOD mode and POD modes, exemplified by figures 6c(i) and 6d(i). The reduction in correlation at the end of the domain is less severe, and most likely reflects the attenuation of the Hann window. Similarly, the correlation coefficient of the second POD mode with the first HPOD mode is consistently high throughout the domain, with small dips in the correlation coefficient occurring at $x/D=1.55$ and $x/D=3.5$, as shown in figure 5b. The spacing between drops in the correlation coefficient between the first HPOD mode and both the first and second POD modes is approximately $1.9D$, suggesting that these modes correspond to the same propagating structure, with a wavelength of approximately $3.8D$.
The fourth and fifth POD modes are the best matches to the second HPOD mode, with correlation coefficients of 0.86 and 0.96, and $R_{k,j}^{2}$ values of 0.68 and 0.84, respectively. The corresponding phase angles are $1.71\pi$ and $1.21\pi$, which are consistent with the expected $\pi/2$ offset. The correlation of the fourth POD mode with the second HPOD mode at the corresponding phase offset, shown in figure 5c, is close to unity between $x/D=1.25$ and $x/D=2.4$, after which it drops before increasing again from approximately $x/D=2.7$. The downstream increase never reaches unity because the fourth POD mode is asymmetric about the centreline, as shown in figure 3d, whereas the second HPOD mode is symmetric, as shown in figures 4c–d. The correlation of the fifth POD mode with the second HPOD mode at the corresponding phase offset, shown in figure 5d, is consistently high across the measurement domain, except towards the edges of the domain, where the Hann window attenuates the intensity of the HPOD modes. This behaviour reflects the greater symmetry of the fifth POD mode, shown in figure 3e, compared to the fourth POD mode, shown in figure 3d.
The sixth and eighth POD modes are the best matches to the third HPOD mode, with correlation coefficients of 0.87 and 0.80, and $R_{k,j}^{2}$ values of 0.75 and 0.60, respectively. The separation of the corresponding phase angles, $1.49\pi$ and $0.92\pi$, is slightly greater than $\pi/2$. The correlation between the sixth POD mode and the third HPOD mode at the corresponding phase offset, shown in figure 5e, remains close to unity across the measurement domain with regularly spaced troughs where the mode values approach zero up to $x/D=3$. Downstream of this point, the correlation remains at unity for a longer streamwise distance before decreasing towards the domain edge. This extended region of high correlation corresponds to the increased size of the structures in the sixth POD mode at this streamwise location, shown in figure 3f. The substantial decrease in correlation towards the downstream end reflects the reduced size of the structures in the HPOD mode, and the subsequent regions of opposing sign which are visible at the end of the domain in figure 4f(i). The correlation between the eighth POD mode and the third HPOD mode at the corresponding phase offset, shown in figure 5f, is consistently high across the measurement domain, with periodic troughs which increase in wavelength downstream, consistent with the structures in the eighth POD mode, shown in figure 3h(i).
The ninth and tenth POD modes are the best matches to the fourth HPOD mode at phase angles of $1.67\pi$ and $1.16\pi$, respectively. The corresponding correlation coefficients are both 0.86, while the $R_{k,j}^{2}$ values are 0.73 and 0.68, respectively. The correlation between the ninth POD mode and the fourth HPOD mode at the corresponding phase offset, shown in figure 5g, is consistently high across the measurement domain, with troughs where the mode values approach zero. Additionally, the correlation drops towards the end of the measurement domain, with the streamwise component approaching zero at the end of the domain and the transverse component changing sign, as shown in figures 3i(i) and 3i(ii), respectively. The correlation between the tenth POD mode and the fourth HPOD mode at the corresponding phase offset, shown in figure 5h, is also consistently high across the measurement domain. The troughs in the correlation are more widely spaced further downstream, consistent with the increasing wavelength of the structures in the tenth POD mode, shown in figure 3j. The correlation also drops towards the ends of the measurement domain, with the transverse component changing sign at the beginning of the domain. The correlation between the streamwise components becomes negative at the end of the domain, reflecting structures of opposite sign appearing in the HPOD modes, due to the spectral leakage visible in figure 4g.
The $R^{2}_{k,\widetilde{(i,j)}}$ values of the $k^{th}$ HPOD mode and the phase average of the $i^{th}$ and $j^{th}$ POD modes are defined as
$R^{2}_{k,\widetilde{(i,j)}}=1-\frac{\sum_{\boldsymbol{x},\phi}{\Big(\psi^{a}_{k}(\boldsymbol{x},\phi+\Delta\phi_{k,i}^{a})-\widetilde{\psi_{i,j}}(\boldsymbol{x},\phi)\Big)}^{2}}{\sum_{\boldsymbol{x},\phi}{\Big(\widetilde{\psi_{i,j}}(\boldsymbol{x},\phi)\Big)}^{2}},$ (46)
[…]
HPOD Mode $k$ POD Modes $i,j$ $R^{2}_{k,\widetilde{(i,j)}}$ $(R^{2}_{k,\widetilde{(i,j)}})^{u}$ $(R^{2}_{k,\widetilde{(i,j)}})^{v}$
1 1,2 0.83 0.79 0.85
2 4,5 0.80 0.78 0.82
3 6,8 0.67 0.53 0.76
4 9,10 0.70 0.61 0.83
@@
Table 4: $R_{k,(i,j)}^{2}$ values of the phase-averaged HPOD modes and the corresponding POD modes. Superscripts u and v refer to the streamwise and transverse components, respectively.
@@
Figure 6: Phase-averaged (a) streamwise and (c) transverse components of the first and second POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the first HPOD mode, adjusted for the phase angle offset $\Delta\phi_{1,1}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 1 for an animated version of this figure.
@@
The streamwise component of the phase-averaged first and second POD modes, shown in figure 6a, agrees well with that of the first HPOD mode at the adjusted phase angle, shown in figure 6b. The main differences in these structures result from the accentuation of propagating structures by HPOD. Specifically, the structures in the POD mode exhibit greater intensity directly behind the sphere, while those further downstream are more pronounced in the HPOD modes. This difference is also evident in the transverse components, shown in figures 6c and 6d for the POD modes and HPOD modes, respectively. Additionally, the shape of the structures in the near wake of the sphere in the POD modes, shown in figure 6a(iv), differs from that in the HPOD mode, shown in figure 6b(iv). This results from the assumed periodicity of the Hilbert transform, which couples the near-wake structures with those further downstream. Despite these differences, it is evident that the first and second POD modes and the first HPOD mode represent the same propagating structure. This structure consists of planar-antisymmetric streamwise fluctuations and alternating positive and negative transverse fluctuations, consistent with wake flapping. The similarity of the third POD mode, shown in figure 3c, to the first and second suggests that the flapping includes an out-of-plane component.
@@
Figure 7: Phase-averaged (a) streamwise and (c) transverse components of the fourth and fifth POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the second HPOD mode, adjusted for the phase angle offset $\Delta\phi_{2,4}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 2 for an animated version of this figure.
@@
While the streamwise component of the phase-averaged fourth and fifth POD modes, shown in figure 7a, resembles the second HPOD mode, shown in figure 7b, the latter exhibits markedly greater symmetry, attributed to HPOD’s inherent emphasis on propagating features. This is most evident when comparing figures 7a(ii) and 7a(vi) to figures 7b(ii) and 7b(vi). In the HPOD case, the V-shaped structures remain symmetric downstream, whereas they become asymmetric in the POD modes. Additionally, the V-shaped structures extend further upstream structures close to the sphere ($x/D<1.5$) in the POD modes, but remain distinct in the HPOD modes. The transverse component of the second HPOD mode, shown in figure 7d, exhibits greater antisymmetry than the transverse components of the phase average of the fourth and fifth POD modes, shown in figure 7c, particularly upstream of $x/D=3$. The asymmetry in the POD modes likely arises because three-dimensionality in the flow is more pronounced in the structures represented by the fourth and fifth POD modes than that represented by the first and second POD modes. In this case, the use of HPOD proves advantageous for isolating the propagating structures in the flow in the presence of asymmetry in the POD modes. The alternating positive and negative streamwise fluctuations in these modes indicates that they correspond to a pulsating motion in the sphere’s wake.
@@
Figure 8: Phase-averaged (a) streamwise and (c) transverse components of the sixth and eighth POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the third HPOD mode, adjusted for the phase angle offset $\Delta\phi_{3,6}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 3 for an animated version of this figure.
@@
The streamwise component of the phase average of the sixth and eighth POD modes, shown in figure 8a, is qualitatively similar to that of the third HPOD mode, shown in figure 8b. However, the alternating positive and negative structures are more distinctly defined in the HPOD mode than the phase average of the POD modes. This is particularly evident when comparing figures 8a(i) and 8a(ii) to figures 8b(i) and 8b(ii), in which the positive region in the HPOD modes around $x/D=3$ and $y/D>0$ is not present in the phase-averaged POD modes. This likely attributable to the out-of-plane motion of the structure as it propagates, while the Hilbert transform enforces periodicity. Additionally, the intensity of the structures tends to be lower in the HPOD modes than those in the POD modes because the non-propagating components are relegated from the leading HPOD modes. The transverse component of the phase average of the sixth and eighth POD modes, shown in figure 8c, exhibits transverse displacement of alternating-sign structures, most evident in figures 8c(i)–(iii). This shifting is absent from the transverse component of the third HPOD mode, shown in figure 8d. The shifting of the structures in the transverse velocity of the POD modes aligns with the three-dimensional motion of the wake and corresponds to the observed behaviour of their streamwise components. The structure represented by these modes suggests wake flapping with a shorter wavelength than that associated with the first and second POD modes and the first HPOD mode.